OpenStudy (anonymous):

what is the derivative of f(x) = cos[sec^-1(e^x)]

OpenStudy (p0sitr0n):

sec^-1 is 1/sec is cos f(x)=cos(cos(e^x) f'(x)= sin(cos(e^x))sin(e^x)e^x

OpenStudy (anonymous):

-e^-x

OpenStudy (anonymous):

So f(x)= sin(cos(e^x))sin(e^x)-e^-x?

OpenStudy (anonymous):

my answer is correct but i m lazy to type everything down...he messed up in chaain rule